A cylinder contains 30.0 L of oxygen gas at a pressure of 1.7 atm and a temperature of 306 K .

How many moles of gas are in the cylinder?
Answer #1

Given:

P = 1.7 atm

V = 30.0 L

T = 306.0 K

find number of moles using:

P * V = n*R*T

1.7 atm * 30 L = n * 0.08206 atm.L/mol.K * 306 K

n = 2.031 mol

Answer: 2.03 mol